Assume that the forecasted demand for 2006 is 15. Use the following date set and exponential smoothing with a(alpha)=0.4 and determine the forecasted demand for 2009. (round to one decimal place)
year 2006. demand 16
year 2007. demand 20
year 2008. demand 18
Answer #1

Smoothing constant = alpha = 0.4

The forecast for period (n + 1) = alpha*(Demand for period n) + (1 - alpha)*Forecast for period n

Fn+1 = alpha*Dn + (1 - alpha)*Fn

Forecast for year 2006 = 15 and Demand = 16

Forecast for year 2007 = 0.4*(16) + (1-0.40)*(15) = 6.4 + 9 = 15.4

Forecast for year 2008 = 0.4*(20) + (1-0.40)*(15.4) = 8 + 9.24 = 17.24

Forecast for year 2009 = 0.4*(18) + (1-0.40)*(17.24) = 7.2 + 10.344 = 17.544